# Acoustic Neural 3D Reconstruction Under Pose Drift

### v1

@@ IV-A Loss Function
We define the following set of trainable parameters: $\Theta$ are the weights of the SDF network $M$, $\Phi$ are the weights of the neural renderer $N$, and $\mathcal{T}=\{T_{i}\}$ which parametrize the learnable sonar poses. Any intensity value computed via Eq. 2 is a function of $\Theta$, $\Phi$, and $\mathcal{T}$ in other words, we can define $\hat{I}_{p}(\Theta,\Phi,\mathcal{T})$ as the predicted intensity of the $p$th pixel of a training image and express our rendering loss function in terms of these three sets of parameters.
@@
$\mathcal{L}_{\text{int}}\equiv\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}||\hat{I}_{p}(\Theta,\Phi,\mathcal{T})-I_{p}||_{1},$ (5)
@@
$\mathcal{L}_{\text{eik}}\equiv\frac{1}{|\mathbf{X}|}\sum_{\mathbf{x}\in\mathbf{X}}(||\nabla\mathbf{N}(\mathbf{x}(\mathcal{T})||_{2}-1)^{2},$ (6)
@@
$\mathcal{L}_{\text{reg}}\equiv\frac{1}{|\mathbf{X}|}\sum_{\mathbf{x}\in\mathbf{X}}||\alpha[\mathbf{x}(\mathcal{T})]||_{1},$ (7)
@@
Since our loss function is a fully differentiable function of all its parameters, we can perform parameter updates using iterative gradient-based methods such as ADAM. After convergence, we retrieve the surface by extracting the zero-level set of $\mathbf{N}$ using the Marching Cubes algorithm:
@@ IV-B Sensor Pose Parametrization
Datasets Real Boat 1 Boat 2 Plane 1 Plane 2 Rock 1 Rock 2 Concrete column Submarine
@@ V Evaluation
We trained our models on an NVIDIA H100 80GB HBM3 GPU with Intel Xeon 8470 CPU. Each training runs for $100$k iterations, which takes about 5 hours. Table I provides the total number of sonar/pose pairs for each simulated and real dataset.
@@ V-C Simulation
NeuSIS (GT) NeuSIS (drift) Ours
RMS Mean       RMS Mean     RMS Mean
@@
    Add noise to the $x,y,\psi$ axes of $\Delta T_{i,i+1}^{\text{dvl}}$ following Eq. 16 with $\varepsilon^{x},\varepsilon^{y}\sim\mathcal{N}(0,0.004$m)))) and $\varepsilon^{\psi}\sim\mathcal{N}(0,0.004\text{rad})$. We obtain the noisy relative transform $\Delta\tilde{T}_{i\rightarrow i+1}^{\text{dvl}}$.
@@ V-D Water Tank Experiments
We evaluate our proposed method on real-world datasets of a test structure submerged in a test tank (see Fig. 4) imaged with two different elevation apertures: $14^{\circ}$ and $28^{\circ}$. Our experimental platform for dataset collections is a Bluefin Hovering Autonomous Underwater Vehicle (HAUV) [44] equipped with a 1.2MHz Teledyne/RDI Workhorse Navigator Doppler velocity log (DVL), a Honeywell HG1700 inertial measurement unit (IMU), a Paroscientific Digiquartz depth sensor, and a Sound Metric DIDSON imaging sonar [45] (please check [10, 5] for more details about the datasets and the collection setup). The frame rate of the real-world datasets is 2Hz.
@@ VI Do we Recover the Odometry Poses?
    Inject noise to relative pose: Corrupt the odometry poses by injecting noise to the $x,y,\psi$ relative poses with $\varepsilon^{x},\varepsilon^{y}\sim\mathcal{N}(0,0.01$m)))) and $\varepsilon^{\psi}\sim\mathcal{N}(0,0.01\text{rad})$. We obtain a noisy drifting set of poses: $\mathcal{T}_{\text{noisy}}$.
@@
Fig. 7 shows the errors between $\mathcal{T}_{\text{odom}}$ and $\mathcal{T}_{\text{noisy}}^{*}$ in the sonar frame, revealing significant differences between the two sets of poses. This indicates that if we freeze the converged mesh $\mathcal{R}_{\text{odom}}$ and optimize only the poses, the optimization can still converge to a solution set far from the odometric poses. Fig. 8 shows the resulting mesh reconstruction when we freeze $\mathcal{T}_{\text{noisy}}^{*}$ and train $\mathbf{M}$ and $\mathbf{N}$ to obtain $\mathcal{R}_{\text{noisy}}^{*}$.
The reconstruction $\mathcal{R}^{*}_{\text{noisy}}$ captures the major parts of the structure (two legs and the middle tile), supporting the observation that even significantly different pose sets can produce plausible 3D reconstructions. This result is supported by the quantitative metrics for $\mathcal{R}_{\text{noisy}}^{*}$ (RMS = $0.077$m, mean = $0.058$m), which are close to the metrics reported when using $\mathcal{T}_{\text{odom}}$ in Table III.
@@ VIII Acknowledgment
T.L., M.Q., and M.K were partially supported by the Office of Naval Research (ONR) grant N00014-24-1-2272 K.Z. and C.A.M. were partially supported by AFOSR award no. FA9550-22-1-0208, NSF award no. 2339616, and ONR award no. N000142312752.

### v2

@@ IV-A Loss Function
We define the following set of trainable parameters: $\Theta$ are the weights of the SDF network $M$, $\Phi$ are the weights of the neural renderer $N$, and $\mathcal{T}=\{T_{i}\}$ which parametrize the learnable sonar poses. Any intensity value computed via Eq. 2 is a function of $\Theta$, $\Phi$, and $\mathcal{T}$ - in other words, we can define $\hat{I}_{p}(\Theta,\Phi,\mathcal{T})$ as the predicted intensity of the $p$th pixel of a training image and express our rendering loss function in terms of these three sets of parameters.
@@
[…]
@@
[…]
@@
[…]
@@
Since our loss function is a fully differentiable function of all its parameters, we can perform parameter updates using iterative gradient-based methods such as ADAM. After convergence, we retrieve the surface by extracting the zero-level set of $\mathbf{N}$ using the Marching Cubes algorithm with a bounding box enclosing the object:
@@ IV-B Sensor Pose Parametrization
[…]
@@ V Evaluation
We trained our models on an NVIDIA H100 80GB HBM3 GPU with Intel Xeon 8470 CPU. Each training runs for 100100k iterations, which takes about 5 hours. Table I provides the total number of sonar/pose pairs for each simulated and real dataset.
@@ V-C Simulation
[…]
@@
    Add noise to the $x,y,\psi$ axes of $\Delta T_{i,i+1}^{\text{dvl}}$ following Eq. 16 with $\varepsilon^{x},\varepsilon^{y}\sim\mathcal{N}(0,0.004$m)) and $\varepsilon^{\psi}\sim\mathcal{N}(0,0.004\text{rad})$. We obtain the noisy relative transform $\Delta\tilde{T}_{i\rightarrow i+1}^{\text{dvl}}$.
@@ V-D Water Tank Experiments
We evaluate our proposed method on real-world datasets of a test structure submerged in a test tank (see Fig. 4) imaged with the two wide elevation apertures achievable by the sonar: $14^{\circ}$ and $28^{\circ}$. Our experimental platform for dataset collections is a Bluefin Hovering Autonomous Underwater Vehicle (HAUV) [44] equipped with a 1.2MHz Teledyne/RDI Workhorse Navigator Doppler velocity log (DVL), a Honeywell HG1700 inertial measurement unit (IMU), a Paroscientific Digiquartz depth sensor, and a Sound Metric DIDSON imaging sonar [45] (please check [10, 5] for more details about the datasets and the collection setup). The frame rate of the real-world datasets is 2Hz.
@@
((a)) 3D reconstruction results for the $14^{\circ}$ elevation aperture real sonar datasets.
((b)) 3D reconstruction results for the $28^{\circ}$ elevation aperture real sonar datasets.
@@ VI Do we Recover the Odometry Poses?
    Inject noise to relative pose: Corrupt the odometry poses by injecting noise to the $x,y,\psi$ relative poses with $\varepsilon^{x},\varepsilon^{y}\sim\mathcal{N}(0,0.01$m)) and $\varepsilon^{\psi}\sim\mathcal{N}(0,0.01\text{rad})$. We obtain a noisy drifting set of poses: $\mathcal{T}_{\text{noisy}}$.
@@
Fig. 7 shows the errors between $\mathcal{T}_{\text{odom}}$ and $\mathcal{T}_{\text{noisy}}^{*}$ in the sonar frame, revealing significant differences between the two sets of poses. This indicates that if we freeze the converged mesh $\mathcal{R}_{\text{odom}}$ and optimize only the poses, the optimization can still converge to a solution set far from the odometric poses. Fig. 8 shows the resulting mesh reconstruction when we freeze $\mathcal{T}_{\text{noisy}}^{*}$ and train $\mathbf{M}$ and $\mathbf{N}$ to obtain $\mathcal{R}_{\text{noisy}}^{*}$. The reconstruction $\mathcal{R}^{*}_{\text{noisy}}$ captures the major parts of the structure (two legs and the middle tile), supporting the observation that even significantly different pose sets can produce plausible 3D reconstructions. This result is supported by the quantitative metrics for $\mathcal{R}_{\text{noisy}}^{*}$ (RMS = 0.0770.077m, mean = 0.0580.058m), which are close to the metrics reported when using $\mathcal{T}_{\text{odom}}$ in Table III.
@@ VIII Acknowledgment
T.L., M.Q., and M.K. were partially supported by the Office of Naval Research (ONR) grant N00014-24-1-2272. K.Z. and C.A.M. were partially supported by AFOSR award no. FA9550-22-1-0208, NSF award no. 2339616, and ONR award no. N000142312752.